# Asymptotic behaviour of the Hitchin metric on the moduli space of Higgs bundles

### v2

@@ 4.3.2 Asymptotically harmonic representatives around the discriminant
We fix a base point of each connected component of $\Sigma_{\theta|U}$. We set $(E_{U},\theta_{U})=(E,\theta)_{|U}$. Note that there exists a non-degenerate symmetric product $C_{U}$ of $(E_{U},\theta_{U})$ compatible with the limiting configuration $h_{\infty}$. Let $h_{U,t}$ be the harmonic metric of $(E_{U},t\theta_{U})$ compatible with $C_{U}$. There exists a holomorphic function $\alpha_{U}$ on $\Sigma_{\theta|U}$ such that $d\alpha_{U}=\nu_{|\Sigma_{\theta|U}}$ and that $\alpha_{U}$ is 00 at the base points of the connected components of $\Sigma_{\theta|U}$. We obtain $F_{\alpha_{U}}\in\mathcal{H}^{0}(\mathop{\rm Def}\nolimits(E_{U},\theta_{U};C_{U}))$ as in §2.2.7. We set
@@
(¯$(\overline{\partial}_{E_{U}}+\mathop{\rm ad}\nolimits t\theta_{U}){\tt H}_{U}(\nu,t)=0,\quad(\overline{\partial}_{E_{U}}+\mathop{\rm ad}\nolimits t\theta_{U})^{\ast}_{h_{U,t},g_{X}}{\tt H}_{U}(\nu,t)=0.$
@@
For any $21/4<r_{2}<r_{1}<$ 2, there exist positive constants $B_{i}>0$ $(i=0,1,2,3)$ such that the following holds on $U(r_{1})\setminus U(r_{2})$ for any $t\geq 1$:
$|F_{\alpha_{U}}|\leq B_{0}\|\nu\|,\quad\quad|{\tt H}_{U}(\nu,t)-F_{\nu}|_{h_{U,t}}\leq B_{1}e^{-B_{2}t}\|\nu\|,\quad\quad|{\tt H}_{U}(\nu,t)|_{h_{U,t}}\leq B_{3}\|\nu\|.$
@@
There exits a unique element $\sigma_{U}(\nu,t)\in A^{0}(U\cap X^{\circ},\mathop{\rm End}\nolimits^{\mathop{\rm asym}\nolimits}(E;C_{U}))$ such that
(¯$(\overline{\partial}_{E_{U}}+\mathop{\rm ad}\nolimits t\theta_{U})\sigma_{U}(\nu,t)={\tt H}_{U}(\nu,t)-F_{\nu}.$
For any $21/4<r_{2}<r_{1}<$ 2, there exist $B_{i}>0$ $(i=1,2)$ such that the following holds on $U(r_{1})\setminus U(r_{2})$ such that
@@
For any $20<r<$ 2, let $\|{\tt H}_{U}(\nu,t)_{|U(r)}\|_{L^{2},h_{U,t}}$ denote the $L^{2}$-norm of ${\tt H}_{U}(\nu,t)$ on $U(r)$ with respect to $h_{U,t}$. Note that it is independent of the choice of a conformal metric of $X$.
@@
For any $20<r<$ 2, there exists $B>0$ such that $\|{\tt H}_{U}(\nu,t)_{|U(r)}\|_{L^{2},h_{U,t}}\leq B\|\nu\|$ for any $t\geq 1$.
@@
For any $20<r<$ 2, there exists $B^{\prime}>0$ such that $|{\tt H}_{U}(\nu,t)|_{h_{U,t},g_{X}}\leq B^{\prime}t\|\nu\|$ on $U(r)$ for any $t\geq 1$.
Proof   We set $Y(t)=\{|w|<2t\}\subset{\mathbb{C}}$. We define the isomorphism $\varphi_{t}:Y(t)\to U$ by $\varphi_{t}(w)=t^{-1}w$. We obtain the harmonic bundles $(\widetilde{E}_{U,t},\widetilde{\theta}_{U,t},\widetilde{h}_{U,t})=\varphi_{t}^{\ast}(E_{U},t\theta_{U},h_{U,t})$ on $Y(t)$. Let $f_{U,t}$ be the endomorphism of $\widetilde{E}_{U,t}$ defined by $\widetilde{\theta}_{U,t}=f_{U,t}\,dw$. Then, there exists $B_{1}>0$ such that any eigenvalue $\alpha$ of $f_{U,t|Q}$ satisfy $|\alpha|\leq B_{1}$ for any $t\geq 1$ and any $Q\in Y(t)$. Let $10<r_{1}<r_{2}<$ 1. By Lemma 4.15, we have
@@
|(¯$\bigl{|}(\overline{\partial}_{E_{U}}+\mathop{\rm ad}\nolimits t\theta_{U})^{\ast}_{h_{t},g_{X}}{\tt H}_{U,t}\bigr{|}_{h_{t}}\leq B_{1}\exp(-B_{2}t)\|\nu\|.$ (42)
Proof   As remarked in Lemma 4.11, we have (¯$(\overline{\partial}_{E_{U}}+\mathop{\rm ad}\nolimits t\theta_{U})^{\ast}_{h_{U,t},g_{X}}{\tt H}_{U,t}=0$. By the estimate $h_{U,t}-h_{t}=O(e^{-\beta t})$ with respect to $h_{U,t}$ for some $\beta>0$ in [23, Theorem 7.17], we obtain (42) from Lemma 4.16.
@@
Proof   By [23, Theorem 7.17], there exists $\beta>0$ such that $h_{U,t}-h_{t}=O(e^{-\beta t})$ on $U(1)$ as $t\to\infty$. Hence, by Lemma 4.15, there exist $B_{i}>0$ $(i=3,4)$ such that the following holds:
@@ 4.3.3 Asymptotically harmonic representatives on $X$
  • (¯$(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta){\tt H}^{\prime}(\nu,t)=0$, and $\rho_{t}({\tt H}^{\prime}(\nu,t))=\pi_{\ast}(\nu)$. In particular, $\mathop{\rm tr}\nolimits({\tt H}^{\prime}(\nu,t))=\mathop{\rm tr}\nolimits\pi_{\ast}(\nu)$. Here, $\rho_{t}$ denotes the morphism $\rho$ (see Lemma 2.11) for $(E,t\theta)$.
@@
  • There exist $B_{i}>0$ $(i=3,4)$ such that |(¯$\bigl{|}(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)_{h_{t},g_{X}}^{\ast}{\tt H}^{\prime}(\nu,t)\bigr{|}_{h_{t}}\leq B_{3}\exp(-B_{4}t)$ on $X$ for any $t\geq 1$.
Proof   We obtain the first two claims by the construction. We obtain the third claim from the estimates for $F_{\nu}$ and ${\tt H}_{\nu}$. The fourth claim follows from Lemma 4.8 and Corollary 4.17.
@@ 4.3.4 Harmonic representatives
(¯$(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)^{\ast}_{h_{t},g_{X}}(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)\gamma(\nu,t)=(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)^{\ast}_{h_{t},g_{X}}{\tt H}^{\prime}(\nu,t).$
@@
Then, ${\tt H}(\nu,t)$ is a harmonic $1$-form satisfying $[{\tt H}(\nu,t)]=[{\tt H}^{\prime}(\nu,t)]=\iota^{{\tt h}}(\nu)$.
@@ 4.4 Harmonic representatives of vertical infinitesimal deformations
Let $\tau$ be a harmonic $(0,1)$-form on $\Sigma_{\theta}$, i.e., $\tau$ is a $(0,1)$-form such that $\partial_{\Sigma_{\theta}}\tau=0$. It induces $[\tau]\in H^{1}(\Sigma_{\theta},\mathcal{O}_{\Sigma_{\theta}})$. Conversely, any element of $H^{1}(\Sigma_{\theta},\mathcal{O}_{\Sigma_{\theta}})$ can be expressed uniquely in this way. We shall study the harmonic representative of ${\tt V}(\tau,t)$ of $\iota_{t}^{{\tt v}}([\tau])$. Let $\|\tau\|$ denote the $L^{2}$-norm of $\tau$ on $\Sigma_{\theta}$.
@@ 4.4.1 Preliminary
In general, for any complex manifold $Y$, let $Y^{\dagger}$ denote the conjugate of $Y$. There exist the natural isomorphism $T^{\ast}(X^{\dagger})\simeq(T^{\ast}X)^{\dagger}$.
@@ 4.4.2 Asymptotically harmonic representatives on $X^{\circ}$
We have $\mathop{\rm tr}\nolimits F_{\tau}=\mathop{\rm tr}\nolimits(\pi^{\dagger}_{1})_{\ast}(\tau)$. In particular, (¯$(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)^{\ast}_{h_{t},g_{X}}\mathop{\rm tr}\nolimits F_{\tau}=0$.
@@ 4.4.3 Asymptotically harmonic $1$-forms around the discriminant
Let $(U,z)$ be as in §4.3.2. We choose a base point of each connected component of $\Sigma_{\theta|U}$. There exists an anti-holomorphic function $\beta_{U}$ such that ¯$\overline{\partial}\beta_{U}=\tau$ on $\Sigma_{\theta|U}$ and that $\beta_{U}=0$ at the base points of the connected components.
@@
(¯$(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta){\tt V}_{U}(\tau,t)=0,\quad\quad(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)^{\ast}_{h_{t},g_{X}}{\tt V}_{U}(\tau,t)=0.$
@@
For any $21/4<r_{2}<r_{1}<$ 2, there exist $B_{i}>0$ $(i=0,1,2,3)$ such that the following holds on $U(r_{1})\setminus U(r_{2})$:
@@
For any $20<r<$ 2, let $\|{\tt V}_{U}(\tau,t)_{|U(r)}\|_{L^{2},h_{t}}$ denote the $L^{2}$-norm of ${\tt V}_{U}(\tau,t)$ on $U(r)$ with respect to $h_{t}$. We obtain the following lemma by using Proposition 2.27 and Lemma 4.24, as in the proof of Lemma 4.15.
@@
For any $20<r<$ 2, there exists $B>0$ such that $\|{\tt V}_{U}(\tau,t)_{|U(r)}\|_{L^{2},h_{t}}\leq B\|\tau\|$ for any $t\geq 1$.
@@
For any $20<r<$ 2, there exists $B^{\prime}>0$ such that $|{\tt V}_{U}(\tau,t)|_{h_{t},g_{X}}\leq B^{\prime}t\|\tau\|$ on $U(r)$ for any $t\geq 1$.
@@ 4.4.4 Asymptotically harmonic representatives on $X$
We obtain a $1$-cocycle $F_{\tau^{\prime}}\in A^{1}(\mathop{\rm Def}\nolimits(E,t\theta))$ such that $[F_{\tau^{\prime}}]=\iota^{{\tt v}}_{t}([\tau])$. We set
@@
  • (¯$(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta){\tt V}^{\prime}(\tau,t)=0$ and $[{\tt V}^{\prime}(\tau,t)]=\iota^{{\tt v}}_{t}([\tau])$.
  • We have $\mathop{\rm tr}\nolimits{\tt V}^{\prime}(\tau,t)=\mathop{\rm tr}\nolimits(\pi^{\dagger}_{1})_{\ast}(\tau)$. In particular, (¯$(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)^{\ast}_{h_{t},g_{X}}\mathop{\rm tr}\nolimits{\tt V}^{\prime}(\tau,t)=0$.
  • There exists $B_{i}>0$ $(i=1,2)$ such that $|{\tt V}^{\prime}(\tau,t)-F_{\tau}|_{h_{t},g_{X}}\leq B_{1}\exp(-B_{2}t)\|\tau\|$ on $X\setminus\bigcup U_{j}(r_{2})$ and that $|{\tt V}^{\prime}(\tau,t)-{\tt V}_{U_{j}}(\tau,t)|_{h_{t},g_{X}}\leq B_{1}\exp(-B_{2}t)\|\tau\|$ on $U_{j}(r_{1})$.
@@
  • There exist $B_{i}>0$ $(i=1,2)$ such that |(¯$\bigl{|}(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)^{\ast}_{h_{t},g_{X}}{\tt V}^{\prime}(\tau,t)\bigr{|}_{h_{t}}\leq B_{1}\exp(-B_{2}t)\|\tau\|$.

### v3

@@ 4.3.2 Asymptotically harmonic representatives around the discriminant
We fix a base point of each connected component of $\Sigma_{\theta|U}$. We set $(E_{U},\theta_{U})=(E,\theta)_{|U}$. Note that there exists a non-degenerate symmetric product $C_{U}$ of $(E_{U},\theta_{U})$ compatible with the limiting configuration $h_{\infty}$. Let $h_{U,t}$ be the harmonic metric of $(E_{U},t\theta_{U})$ compatible with $C_{U}$. There exists a holomorphic function $\alpha_{U}$ on $\Sigma_{\theta|U}$ such that $d\alpha_{U}=\nu_{|\Sigma_{\theta|U}}$ and that $\alpha_{U}$ is 0 at the base points of the connected components of $\Sigma_{\theta|U}$. We obtain $F_{\alpha_{U}}\in\mathcal{H}^{0}(\mathop{\rm Def}\nolimits(E_{U},\theta_{U};C_{U}))$ as in §2.2.7. We set
@@
$(\overline{\partial}_{E_{U}}+\mathop{\rm ad}\nolimits t\theta_{U}){\tt H}_{U}(\nu,t)=0,\quad(\overline{\partial}_{E_{U}}+\mathop{\rm ad}\nolimits t\theta_{U})^{\ast}_{h_{U,t},g_{X}}{\tt H}_{U}(\nu,t)=0.$
@@
For any $1/4<r_{2}<r_{1}<2$, there exist positive constants $B_{i}>0$ $(i=0,1,2,3)$ such that the following holds on $U(r_{1})\setminus U(r_{2})$ for any $t\geq 1$:
$|F_{\alpha_{U}}|_{h_{U,t}}\leq B_{0}\|\nu\|,\quad\quad|{\tt H}_{U}(\nu,t)-F_{\nu}|_{h_{U,t}}\leq B_{1}e^{-B_{2}t}\|\nu\|,\quad\quad|{\tt H}_{U}(\nu,t)|_{h_{U,t}}\leq B_{3}\|\nu\|.$
@@
There exists a unique element $\sigma_{U}(\nu,t)\in A^{0}(U\cap X^{\circ},\mathop{\rm End}\nolimits^{\mathop{\rm asym}\nolimits}(E;C_{U}))$ such that
$(\overline{\partial}_{E_{U}}+\mathop{\rm ad}\nolimits t\theta_{U})\sigma_{U}(\nu,t)={\tt H}_{U}(\nu,t)-F_{\nu}.$
For any $1/4<r_{2}<r_{1}<2$, there exist $B_{i}>0$ $(i=1,2)$ such that the following holds on $U(r_{1})\setminus U(r_{2})$ such that
@@
For any $0<r<2$, let $\|{\tt H}_{U}(\nu,t)_{|U(r)}\|_{L^{2},h_{U,t}}$ denote the $L^{2}$-norm of ${\tt H}_{U}(\nu,t)$ on $U(r)$ with respect to $h_{U,t}$. Note that it is independent of the choice of a conformal metric of the Riemann surface $X$.
@@
For any $0<r<2$, there exists $B>0$ such that $\|{\tt H}_{U}(\nu,t)_{|U(r)}\|_{L^{2},h_{U,t}}\leq B\|\nu\|$ for any $t\geq 1$.
@@
For any $0<r<2$, there exists $B^{\prime}>0$ such that $|{\tt H}_{U}(\nu,t)|_{h_{U,t},g_{X}}\leq B^{\prime}t\|\nu\|$ on $U(r)$ for any $t\geq 1$.
Proof  We set $Y(t)=\{|w|<2t\}\subset{\mathbb{C}}$. We define the isomorphism $\varphi_{t}:Y(t)\to U$ by $\varphi_{t}(w)=t^{-1}w$. We obtain the harmonic bundles $(\widetilde{E}_{U,t},\widetilde{\theta}_{U,t},\widetilde{h}_{U,t})=\varphi_{t}^{\ast}(E_{U},t\theta_{U},h_{U,t})$ on $Y(t)$. Let $f_{U,t}$ be the endomorphism of $\widetilde{E}_{U,t}$ defined by $\widetilde{\theta}_{U,t}=f_{U,t}\,dw$. Then, there exists $B_{1}>0$ such that any eigenvalue $\alpha$ of $f_{U,t|Q}$ satisfies $|\alpha|\leq B_{1}$ for any $t\geq 1$ and any $Q\in Y(t)$. Let $0<r_{1}<r_{2}<1$. By Lemma 4.15, we have
@@
$\bigl|(\overline{\partial}_{E_{U}}+\mathop{\rm ad}\nolimits t\theta_{U})^{\ast}_{h_{t},g_{X}}{\tt H}_{U,t}\bigr|_{h_{t}}\leq B_{1}\exp(-B_{2}t)\|\nu\|.$ (46)
Proof  As remarked in Lemma 4.11, we have $(\overline{\partial}_{E_{U}}+\mathop{\rm ad}\nolimits t\theta_{U})^{\ast}_{h_{U,t},g_{X}}{\tt H}_{U,t}=0$. By the estimate $h_{U,t}-h_{t}=O(e^{-\beta t})$ with respect to $h_{U,t}$ for some $\beta>0$ in [24, Theorem 7.17], we obtain (46) from Lemma 4.16.
@@
Proof  By [24, Theorem 7.17], there exists $\beta>0$ such that $h_{U,t}-h_{t}=O(e^{-\beta t})$ on $U(1)$ as $t\to\infty$. Hence, by Lemma 4.15, there exist $B_{i}>0$ $(i=3,4)$ such that the following holds:
@@ 4.3.3 Asymptotically harmonic representatives on $X$
  • $(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta){\tt H}^{\prime}(\nu,t)=0$, and $\rho_{t}({\tt H}^{\prime}(\nu,t))=\pi_{\ast}(\nu)$. In particular, $\mathop{\rm tr}\nolimits({\tt H}^{\prime}(\nu,t))=\mathop{\rm tr}\nolimits\pi_{\ast}(\nu)$. Here, $\rho_{t}$ denotes the morphism $\rho$ (see Lemma 2.11) for $(E,t\theta)$.
@@
  • There exist $B_{i}>0$ $(i=3,4)$ such that $\bigl|(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)_{h_{t},g_{X}}^{\ast}{\tt H}^{\prime}(\nu,t)\bigr|_{h_{t}}\leq B_{3}\exp(-B_{4}t)$ on $X$ for any $t\geq 1$.
  • $H^{\prime}(\nu,t)$ is horizontal, i.e., $H^{\prime}(\nu,t)]=\iota^{{\tt h}}_{t}(\nu)$.
Proof  We obtain the first claim by the construction. We obtain the second claim by the construction and the estimate in Lemma 4.13. We obtain the third claim from the estimates for $F_{\nu}$ and ${\tt H}_{\nu}$. The fourth claim follows from Lemma 4.8 and Corollary 4.17. We obtain the last claim by Corollary 3.32 and the construction of $H^{\prime}(\nu,t)$.
@@ 4.3.4 Harmonic representatives
$(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)^{\ast}_{h_{t},g_{X}}(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)\gamma(\nu,t)=(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)^{\ast}_{h_{t},g_{X}}{\tt H}^{\prime}(\nu,t).$
@@
Then, ${\tt H}(\nu,t)$ is a harmonic $1$-form satisfying ${\tt H}(\nu,t)]=[{\tt H}^{\prime}(\nu,t)]=\iota^{{\tt h}}(\nu)$.
@@ 4.4 Harmonic representatives of vertical infinitesimal deformations
Let $\tau$ be a harmonic $(0,1)$-form on $\Sigma_{\theta}$, i.e., $\tau$ is a $(0,1)$-form such that $\partial_{\Sigma_{\theta}}\tau=0$. It induces $\tau]\in H^{1}(\Sigma_{\theta},\mathcal{O}_{\Sigma_{\theta}})$. Conversely, any element of $H^{1}(\Sigma_{\theta},\mathcal{O}_{\Sigma_{\theta}})$ can be expressed uniquely in this way. We shall study the harmonic representative of ${\tt V}(\tau,t)$ of $\iota_{t}^{{\tt v}}([\tau])$. Let $\|\tau\|$ denote the $L^{2}$-norm of $\tau$ on $\Sigma_{\theta}$.
@@ 4.4.1 Preliminary
In general, for any complex manifold $Y$, let $Y^{\dagger}$ denote the conjugate of $Y$. There exists the natural isomorphism $T^{\ast}(X^{\dagger})\simeq(T^{\ast}X)^{\dagger}$.
@@ 4.4.2 Asymptotically harmonic representatives on $X^{\circ}$
We have $\mathop{\rm tr}\nolimits F_{\tau}=\mathop{\rm tr}\nolimits(\pi^{\dagger}_{1})_{\ast}(\tau)$. In particular, $(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)^{\ast}_{h_{t},g_{X}}\mathop{\rm tr}\nolimits F_{\tau}=0$.
@@ 4.4.3 Asymptotically harmonic $1$-forms around the discriminant
Let $(U,z)$ be as in §4.3.2. We choose a base point of each connected component of $\Sigma_{\theta|U}$. There exists an anti-holomorphic function $\beta_{U}$ such that $\overline{\partial}\beta_{U}=\tau$ on $\Sigma_{\theta|U}$ and that $\beta_{U}=0$ at the base points of the connected components.
@@
$(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta){\tt V}_{U}(\tau,t)=0,\quad\quad(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)^{\ast}_{h_{t},g_{X}}{\tt V}_{U}(\tau,t)=0.$
@@
For any $1/4<r_{2}<r_{1}<2$, there exist $B_{i}>0$ $(i=0,1,2,3)$ such that the following holds on $U(r_{1})\setminus U(r_{2})$:
@@
For any $0<r<2$, let $\|{\tt V}_{U}(\tau,t)_{|U(r)}\|_{L^{2},h_{t}}$ denote the $L^{2}$-norm of ${\tt V}_{U}(\tau,t)$ on $U(r)$ with respect to $h_{t}$. We obtain the following lemma by using Proposition 2.27 and Lemma 4.24, as in the proof of Lemma 4.15.
@@
For any $0<r<2$, there exists $B>0$ such that $\|{\tt V}_{U}(\tau,t)_{|U(r)}\|_{L^{2},h_{t}}\leq B\|\tau\|$ for any $t\geq 1$.
@@
For any $0<r<2$, there exists $B^{\prime}>0$ such that $|{\tt V}_{U}(\tau,t)|_{h_{t},g_{X}}\leq B^{\prime}t\|\tau\|$ on $U(r)$ for any $t\geq 1$.
@@ 4.4.4 Asymptotically harmonic representatives on $X$
We obtain a $1$-cocycle $F_{\tau^{\prime}}\in A^{1}(\mathop{\rm Def}\nolimits(E,t\theta))$ such that $F_{\tau^{\prime}}]=\iota^{{\tt v}}_{t}([\tau])$. We set
@@
  • $(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta){\tt V}^{\prime}(\tau,t)=0$ and ${\tt V}^{\prime}(\tau,t)]=\iota^{{\tt v}}_{t}([\tau])$.
  • We have $\mathop{\rm tr}\nolimits{\tt V}^{\prime}(\tau,t)=\mathop{\rm tr}\nolimits(\pi^{\dagger}_{1})_{\ast}(\tau)$. In particular, $(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)^{\ast}_{h_{t},g_{X}}\mathop{\rm tr}\nolimits{\tt V}^{\prime}(\tau,t)=0$.
  • There exist $B_{i}>0$ $(i=1,2)$ such that $|{\tt V}^{\prime}(\tau,t)-F_{\tau}|_{h_{t},g_{X}}\leq B_{1}\exp(-B_{2}t)\|\tau\|$ on $X\setminus\bigcup U_{j}(r_{2})$ and that $|{\tt V}^{\prime}(\tau,t)-{\tt V}_{U_{j}}(\tau,t)|_{h_{t},g_{X}}\leq B_{1}\exp(-B_{2}t)\|\tau\|$ on $U_{j}(r_{1})$.
@@
  • There exist $B_{i}>0$ $(i=1,2)$ such that $\bigl|(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)^{\ast}_{h_{t},g_{X}}{\tt V}^{\prime}(\tau,t)\bigr|_{h_{t}}\leq B_{1}\exp(-B_{2}t)\|\tau\|$.